1/2

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,
Varulven  The Banshee (An Approach)  Der Werwolf
tradukita de Sven Collberg  tradukita de Max Knight

En varulv lopp från fru m.m. en vacker afton samt begav sig till en bymagisters grav och bad: “Var snäll och konjugera.

One night, a banshee slunk away

from mate and child, and in the gloom went to a village teacher’s tomb, requesting him: “Inflect me, pray.”

Ein Werwolf eines Nachts entwich von Weib und Kind und sich begab an eines Dorfschullehrers Grab und bat ihn: �Bitte, beuge mich!“

mig!” Byskolläraren stod upp på blecknamnsskyltens mässingknopp med svar till den, som satt på pass med tåligt korslagd tass på tass.

The village teacher climbed up straight upon his grave stone with its plate and to the apparition said who meekly knelt before the dead:

Der Dorfschulmeister stieg hinauf auf seines Blechschilds Messingknauf und sprach zum Wolf, der seine Pfoten geduldig kreuzte vor dem Toten:

“Jag var-ulv 1:a pers. i sing., du, han, hon, den, det var-ulv. Så vi voro-ulv, I voren. På de voro-ulv kom ingenting.”

“The banSHEE, in the subject’s place; the banHERS, the possessive case. The banHER, next, is what they call objective case–and that is all.”

�Der Werwolf“ – sprach der gute Mann, �des Werwolfs, Genitiv sodann, dem Werwolf, Dativ, wie man’s nennt, den Werwolf, - damit hat’s ein End!“

Varulven myste gladåt hur pers. hade följts av sing. och plur. “Fast”, bad han, “kan det inte ges ens en enda pers. i tempus presens?”

The banshee marveled at the cases

and writhed with pleasure, making faces, but said: “You did not add, so far, the plural to the singular!”

Dem Werwolf schmeichelten die Fälle er rollte seine Augenbälle. �Indessen“, bat er, �füge doch zur Einzahl auch die Mehrzahl noch!“

Bekänna måste vår magister, att därvidlag fanns vissa brister. En varulv sine imperfecto var contradictio in adjecto.

The teacher, though, admitted then that this was not within his ken. “While bans are frequent”, he advised, “A she cannot be plurized.”

Der Dorfschullehrer aber musste

gestehn, dass er von ihr nichts wusste. Zwar Wölfe gab’s in großer Schar, doch �Wer“ gab’s nur im Sinular.

…  …  …

2/2

Varulven satt med tårad lins. Han visste att han fanns och finns. Därtill kom också fru m.m. som även måste existera.

The banshee, rising clammily, wailed: “What about my family?” Then, being not a learned creature, said humbly “Thanks” and left the teacher.

Der Wolf erhob sich tränenblind – er hatte ja doch Weib und Kind! Doch da er keine Gelehrter eben, so schied er dankend und ergeben.

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Svedan de Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Verkinto de tiu ĉi Germana poemo estas Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

http://www.poezio.net/version?poem-id=825&version-id=1654,1660,1652